# Monte Carlo Physics-informed neural networks for multiscale heat conduction via phonon Boltzmann transport equation

### v3

@@ 2.2.2 Two-step sampling approach
To address the issues mentioned in the above two sampling methods, we propose the following two-step sampling strategy at each training step, i.e., (1) we first randomly draw $B_{t,\bm{x}}$ points in the temporal-spatial domain, i.e., $(t,\bm{x})_{i},~{}i=1,...,B_{t,\bm{x}}$ (Step I); and (2) we then randomly draw $B_{\bm{s}}$ points in the solid angular domain, i.e., $\bm{s}_{k},~{}k=1,...,B_{\bm{s}}$ (Step II). The training points in the temporal-spatial-angular space $(t,\bm{x},\bm{s})$ at each training step are constructed using the tensor product given the points sampled at the two steps above. In general, we set $B_{\bm{s}}>1$ to improve the accuracy for the computations of the integral in Eq. (2) as well as $e^{eq}$. Hence, we expect that the proposed two-step sampling approach is able to achieve better accuracy when compared to the second sampling method mentioned in Sec. 2.2.1. Note that for steady problems, we can just drop the dimension $t$, and perform the two-step sampling approach only in the spatial and angular domains, which could improve the computational efficiency.
@@
Here we further discuss the superiority of the two-step sampling approach over the the first sampling method in Sec. 2.2.1. As aforementioned, a large number of points in the solid angular domain has to be employed in the ballistic regime for the first approach of Sec. 2.2.1. In addition, how to determine the optimal number of discrete points at different regimes remains an open question. In the two-step sampling method, we randomly generate a certain number of points in the solid angular space at each training step. We can then traverse the entire angular domain even if we employ a small number of points in each step as long as we have enough training steps, which is similar to the statistic or particle methods for solving the BTE. Hence, compared to the discrete ordinate/velocity methods, the present solver is free of ray effects in the ballistic regime. Therefore, less computational and memory cost for heat conduction at the ballistic regimes are then expected at each training step if the two-step sampling strategy is adopted. Also, as we will show in Sec. 3.1, the two-step sampling method is more flexible than the first sampling approach in Sec. 2.2.1 for multiscale heat conduction problems. For the implementations of boundary/initial conditions, we can randomly generate a certain number of training points in the temporal-spatial-angular domain, i.e., $B_{t,\bm{x},\bm{s}}$, since we do not need to compute the integral in boundary/initial conditions, for computational efficiency.
For the random sampling in solid angular space, as previously discussed, the group velocity of the distribution function in our phonon BTE model is uniformly distributed over the surface of a unit sphere. The points can then be expressed in the Cartesian coordinates as $(\cos\theta,\cos\theta\sin\phi,\sin\theta\sin\phi)$. In the present work, we first randomly sample the projection of the group velocity $\cos\theta\in[-1,1]$ along the $x$-axis, and then sample the azimuthal angle $\phi\in[0,2\pi]$ to generate the random points in the solid angular space.
@@ 3 Results and Discussion
In this section, we test the proposed method for simulating the heat conduction in solids ranging from diffusive to ballistic regimes, using four demonstration examples: (1) steady and unsteady quasi-one-dimensional (quasi-1D) film heat conduction, (2) quasi-two-dimensional (quasi-2D) heat conduction in a square domain, and (3) three-dimensional (3D) cuboid phonon transport. Further, the Adam optimizer, which is a variant of the stochastic gradient descent method, is utilized to minimize the loss functions in MC-PINNs. The details for computations, e.g., the architectures of NNs, activation functions, and the numbers of training steps, etc., as well as the details for reference solutions, are present in A.
@@ 3.1 Quasi-1D film heat conduction
We first consider a quasi-1D heat conduction problem across a film, as depicted in Fig. 3. The thickness of the film is $L=1$, which serves as the characteristic length of the problem. The temperatures at the left and right boundaries, i.e., $x=0$ and $L$, are $T_{h}=1$ and $T_{c}=0$, respectively. Specifically, we employ the MC-PINNs to model the heat conduction at four different Knudsen numbers, i.e., $\rm{Kn}=0.01,0.1,1$, and $10$, to justify the capability of the present method for multiscale heat conduction. In all test cases, we randomly select $B_{\bm{x}}=40$ and $B_{\bm{s}}=16$ points in the spatial and angular domain for the computations of the loss for the equation at each training step. For the quasi-1D problem, the angular variable can be expressed as $s_{x}=\cos{\theta}$. For the boundary conditions, we randomly select $B_{x_{0},\bm{s}}=50$ points in the solid angular space at the boundary $x=0$ and $x=1$ and apply the thermalization boundary condition in Eq. (5).
To justify the accuracy of the present method, we first present the energy distribution functions for the case with $\rm{Kn}=1$ at two representative locations, i.e., $x=0$ and $0.5$, computed from the MC-PINNs in Fig. 4. As shown, the results from the MC-PINNs agree well with numerical results in  [19]. In addition, the temperature as well as the heat flux are quantities of interest (QoI) in heat conduction problems rather than the energy distribution functions. We then illustrate the predicted temperature and the heat flux for $x\in[0,1]$ from MC-PINNs in Fig. 5. It is observed that: (1) all the results from the MC-PINNs are in good agreement with those in [16], (2) the temperature exhibits significant nonlinearities near the boundaries for the test cases with $\rm{Kn}>0.01$ (Fig. 5), and (3) the heat flux $q$ remains constant although the temperature is nonlinear for $\rm{Kn}>0.01$, which is clearly different from the widely used Fourier’s law (Fig. 5). In other words, the heat conduction for cases with $\rm{Kn}>0.01$ exhibits significant non-Fourier behavior.
@@
As shown, the relative $L_{2}$ errors between the present method and the reference solutions for all test cases are less than $1\%$, demonstrating that the MC-PINNs is capable of simulating heat conduction at both diffusive and ballistic scales with good accuracy.
Figure 5: Quasi-1D film heat conduction: Predicted (a) temperature, and (b) heat flux, from MC-PINNs at different Knudsen numbers. Solid line: MC-PINNs; Symbols: Reference solution [16]. Gray dashed line: the solution of Fourier’s Law.
Table 1: Quasi-1D film heat conduction: Relative $L_{2}$ errors between PINNs and reference solution at different $\rm{Kn}$. $E_{\phi}$: Relative $L_{2}$ errors for temperature/heat flux.
@@
    PINNs-2: We employ $B_{\bm{x}}$ and $B_{\bm{s}}$ uniform discrete points in the spatial and solid angular domains, respectively.
@@
    PINNs-3: In the spatial domain, we use $B_{\bm{x}}$ uniform discrete points, and $B_{\bm{x}}$ quadrature points from the Gauss-Legendre method are employed in the solid angular domain, The residual points at each training step is then constructed using the tensor product. We note that this sampling approach is the same as in [27].
@@
    PINNs-4: We randomly sample $B_{\bm{s}}$ points in the solid angular domain, but use $B_{\bm{x}}$ uniform points in the spatial domain. Note that the sampling approach here can be viewed as a special case in PINN-1 by replacing the random points with the fixed uniform grids in the spatial domain.
@@
    PINNs-5: Randomly sample $B_{\bm{x}}\times B_{\bm{s}}$ residual points in the spatial-angular domain at each training step. We note that this approach is one of the most widely used methods in PINNs.
Figure 6: Quasi-1D film heat conduction: Effect of different sampling approaches in solid angular space on the predicted accuracy at (a) $\rm{Kn}=0.01$, (b) $\rm{Kn}=1$, and (c) $\rm{Kn}=10$. Symbols: Reference solution; PINNs-1: Two-step sampling method: Random sampling in both spatial and solid angular domain; PINNs-2: uniform discrete points in both spatial and solid angular domain; PINNs-3: uniform discrete points in spatial domain and GL quadrature points in angular domain; PINNs-4: Two-step sampling method: uniform discrete points in spatial domain and random sampling in solid angular space; PINNs-5: random sampling in spatial-angular domain.
We utilize these five sampling methods in PINNs to model the heat conduction in the 1D film at $\rm{Kn}=0.01$, 1, and 10. The training step is set as 50,000 at $\rm{Kn}=1,10$ and 10,$0$ at $\rm{Kn}=0.01$ in all test cases. The results and the computational errors are present in Fig. 6 and Table 2, respectively. As we can see, (1) PINNs with the two-step sampling approach, i.e., PINNs-1 and PINN-4, are able to provide the most accurate results for the heat conduction from diffusive to ballistic regimes; (2) the method in [27], i.e., PINNs-3, is able to achieve good accuracy for Kn $=0.01$ and 1, but failed to capture the non-Fourier’s behavior accurately at Kn $=10$ (e.g., the temperature jumps at the boundaries), as shown in Fig. 6; and (3) the computational errors for PINN-5 increases with the decreasing Kn. As discussed in Sec. 2.2.2, we only solve the BTE at one point in the solid angular space at each spatial point, leading to poor accuracy for the computation of $e^{eq}$. At the ballistic scale, i.e., cases with large Kn, we have very few collisions for the phonon, and the inaccurate estimation of $e^{eq}$ has little effect on the predicted accuracy. However, for cases with small Kn, we have frequent collisions of the phonon, and the inaccurate prediction of $e^{eq}$ results in poor accuracy for solving the BTE.
Table 2: Quasi-1D film heat conduction: Relative $L_{2}$ errors of the temperatures between PINNs and reference solution with different training points. PINN-1 and PINN-4: Two-step sampling approach proposed in the current study; PINN-3: sampling approach used in [25, 27]; PINN-5: sampling approach used in [26].
@@
$E_{T}(\rm{Kn}=0.01)$ 0.402% 0.383% 0.370% 0.390% 50.10%
$E_{T}(\rm{Kn}=1)$ 0.299% 1.005% 0.678% 0.214% 9.201%
$E_{T}(\rm{Kn}=10)$ 0.106% 1.280% 1.076% 0.130% 0.476%
3.2 Unsteady quasi-1D film heat conduction with varying Knudsen number
In this section, we apply our method to model the time-dependent heat conduction in a 1D film with a varying Knudsen number, which is defined as follows:
@@ 3.2 Unsteady quasi-1D film heat conduction with varying Knudsen number
where $\rm{Kn}_{\max}$ and $\rm{Kn}_{\min}$ represent the maximum and minimum Knudsen number, respectively; $x_{c}=L/2$ is the center of the film with $L$ denoting the thickness of the film, and $d=0.01$ characterizes the thickness of the transition layer. In the present case, we set $\rm{Kn}_{\max}=10$ and $\rm{Kn}_{\min}=0.01$. The distribution of the Knudsen number within the film is shown in Fig. 7. It can be seen that we have a significant change in $\rm{Kn}$ near the center of the film. The heat conduction at the left half is at the ballistic regime, which is at the diffusive regime on the right half of the film.
@@
The computational domain as well as the boundary conditions are kept the same as in Sec. 3.1. In addition, we set $T(t=0,\bm{x})=(T_{h}+T_{c})/2$ in the entire domain. The time domain here is set as $t\in[0,10]$. In the computations, we sample $B_{t,x}=100$ random points in the temporal-spatial domain and $B_{s}=16$ random points in the solid angular domain. Further, $B_{t,x_{0},\bm{s}}=100$ random points are sampled for boundary conditions and $B_{t_{0},x,\bm{s}}=100$ random points for initial conditions.
@@
We would like to point out that 100 Gauss-Legendre quadrature points are employed in the solid angular space in the numerical method to accurately capture the non-Fourier’s behavior [19]. In MC-PINNs, we only use 16 points in the solid angular domain at each training step, and it is capable of providing good results for heat conduction ranging from diffusive to ballistic scales. The above results indicate the great flexibility of the MC-PINNs with the two-step sampling approach. We do not need to carefully design the discretization of the angular domain, which is a long-standing challenge in conventional numerical methods for simulating multiscale heat conduction problems.
Figure 8: Unsteady quasi-1D film heat conduction: The predicted temperature at different times with space varying Knudsen number. (a) $t=0.2$, (b) $t=0.5$, (c) $t=1.0$, (d) $t=10.0$. Symbols: Reference solution [19]; Black solid line: MC-PINNs.
@@ 3.3 Quasi-2D square heat conduction
We now consider a quasi-2D heat conduction problem in a square domain with the length $L=1$. The temperature is set to $T_{h}=1$ at the top boundary and $T_{c}=0$ at the remaining boundaries, as shown in Fig. 9. For the quasi-2D problem, the angular variable $\bm{s}=(s_{x},s_{y})$ can be expressed as follows:
$s_{x}=\cos{\theta},~{}s_{y}=\sin{\theta}\cos{\phi},$ (11)
where $\theta\in[0,\pi]$ and $\phi\in[0,\pi]$.
Figure 9: Schematic of the quasi-2D heat conduction in a square domain. $L$ is the characteristic length, $T_{h}$ is the upper plate non-dimensional temperature and $T_{c}$ is the non-dimensional temperature of the remaining plates.
In this section, we employ the MC-PINNs to model the heat conduction in the 2D square domain at $\rm{Kn}=0.1,1$ and $10$. In our computations, we first randomly sample $B_{\bm{x}}=100$ points in the spatial domain and then select $B_{\bm{s}}=64$ random points in the solid angular domain. As for boundary conditions, we randomly select $B_{\bm{x,s}}=100$ in the spatial-angular domain at each boundary to implement the thermalization boundary condition defined in Eq. (5).
The results of MC-PINNs are shown in Figs. 10 and 11. We can see that the predicted temperatures from the MC-PINNs are in good agreement with the reference solutions in all test cases. We also observe in Figs. 10 and 11 there are significant temperature jumps at the left/right boundaries. In addition, the temperature jump becomes more pronounced with the increase of $\rm{Kn}$. This phenomenon is reasonable since the non-equilibrium effects become more pronounced in the transition and ballistic regimes, leading to distinct temperature discontinuities at the walls.
Figure 10: Quasi-2D heat conduction: Distribution of the temperature at (a) $\rm{Kn}=0.1$, (b) $\rm{Kn}=1.0$ and (c) $\rm{Kn}=10.0$. Colored background with black solid line: reference solution [17]; Red dashed line: the MC-PINNs.
Figure 11: Quasi-2D heat conduction: Temperature along the center vertical line at different $\rm{Kn}$. Solid line: MC-PINNs; Symbols: Reference solution [17].
We further compute the relative $L_{2}$ error between the MC-PINNs and the reference solutions at different $\rm{Kn}$ to demonstrate the accuracy of the present method in quasi-2D problems. As shown in Table 3, the relative $L_{2}$ errors for the three test cases are smaller than $2\%$, which again confirms the good accuracy of MC-PINNs for modeling multiscale heat conduction.
Table 3: Quasi-2D heat conduction problem: Relative $L_{2}$ errors between PINNs and reference solution at different $\rm{Kn}$.
- $\rm{Kn}=0.1$ $\rm{Kn}=1.0$ $\rm{Kn}=10.0$
@@
In this section, we proceed to consider the phonon transport in a 3D cuboid. The temperature is set as $T_{h}$ at the top boundary and $T_{c}$ at the remaining boundaries, as shown in Fig. 12. Specifically, we set $T_{h}=1$, $T_{c}=0$, and $L=1$. For the 3D problem, we have $\bm{s}=(s_{x},s_{y},s_{z})$. Since the magnitude of the direction vector $|\bm{s}|=1$, $\bm{s}$ can then be expressed by $\theta$ and $\phi$, i.e.,
$s_{x}$ $=\cos{\theta},$ (12)
$s_{y}$ $=\sin{\theta}\cos{\phi},$
$s_{z}$ $=\sin{\theta}\sin{\phi},$
where $\theta\in[0,\pi]$ and $\phi\in[0,2\pi]$. In addition, we test the heat conduction at three representative Kn, i.e., Kn $=0.1$, 1, and 10.
@@
In the MC-PINNs, the two-step sampling strategy is applied for generating the training points for the loss of residual with $B_{\bm{x}}=100$ and $B_{\bm{s}}=64$. For the boundary conditions, we randomly select $B_{\bm{x,s}}=100$ in the spatial-angular domain for each wall at each training step. The boundary condition implemented here is defined in Eq. (5), which is the same as the previous two cases in Secs. 3.1 and 3.3. As for the reference solution, we employ the implicit kinetic scheme to solve the stationary phonon BTE developed in [17]. We note that the numerical method in [17] employs a macroscopic equation to accelerate convergence in the diffusive regime. Additionally, it incorporates a memory reduction technique to reduce memory usage in computations. In numerical method the grid size is set to $40^{3}\times 40^{2}$ at $\rm{Kn}=0.1,1$ and $40^{3}\times 80^{2}$ at $\rm{Kn}=10$ to ensure accuracy.
The predicted temperature from MC-PINNs in each case at $x=0.5$ are present in Figs. 13 and 14. Again, good agreements are observed for all test cases between the MC-PINNs and the reference solutions. Similar to in the 2D cases, there are significant temperature jumps at the boundaries when $\rm{Kn}\geq 1$, which have been discussed in Sec. 3.3 and will not present here for simplicity. We then illustrate the relative $L_{2}$ errors between the temperature from the MC-PINNs and the reference solutions at different $\rm{Kn}$ in Table 4. The results further confirm the capability of the MC-PINNs for 3D multiscale heat conduction problems.
@@
Figure 14: 3D heat conduction problem: Predicted temperature from MC-PINNs at $x=0.5,y=0.5$. Solid line: MC-PINNs; Symbols: Reference solution [17].

### v4

@@ 2.2.2 Two-step sampling approach
To address the issues mentioned in the above two sampling methods, we propose the following two-step sampling strategy at each training step, i.e., (1) we first randomly draw $B_{t,\bm{x}}$ points in the temporal-spatial domain, i.e., $(t,\bm{x})_{i},~{}i=1,...,B_{t,\bm{x}}$ (Step I); and (2) we then randomly draw $B_{\bm{s}}$ points in the solid angular domain, i.e., $\bm{s}_{j},~{}j=1,...,B_{\bm{s}}$ (Step II). The training points in the temporal-spatial-angular space $(t,\bm{x},\bm{s})$ at each training step are constructed using the tensor product given the points sampled at the two steps above. In general, we set $B_{\bm{s}}>1$ to improve the accuracy for the computations of the integral in Eq. (3) as well as $e^{eq}$. Hence, we expect that the proposed two-step sampling approach is able to achieve better accuracy when compared to the second sampling method mentioned in Sec. 2.2.1. Note that for steady problems, we can just drop the dimension $t$, and perform the two-step sampling approach only in the spatial and angular domains. In our computations, to obtain the samples of $\bm{s}$, we first randomly sample the projection of the polar angle $\cos\theta\in[-1,1]$ along the $x$-axis, and sample the azimuthal angle $\varphi\in[0,2\pi]$ to generate the random points in the solid angular space.
For the implementations of boundary/initial conditions, we randomly generate a certain number of training points in the temporal-spatial-angular domain, i.e., $B_{t,\bm{x},\bm{s}}$, since we do not need to compute the integral in boundary/initial conditions, for computational efficiency.
@@
Here we further discuss the superiority of the two-step sampling approach over the the first sampling method in Sec. 2.2.1. As aforementioned, a large number of points in the solid angular domain have to be employed in the ballistic regime for the first approach of Sec. 2.2.1. In addition, how to determine the optimal number of discrete points at different regimes remains an open question. In the two-step sampling method, we randomly generate a certain number of points in the solid angular space at each training step. We can then traverse the entire angular domain even if we employ a small number of points in each step as long as we have enough training steps, which is similar to the statistic or particle methods for solving the BTE. Hence, compared to the discrete ordinate/velocity methods, the present solver is free of ray effects in the ballistic regime. Therefore, less computational and memory cost for heat conduction at the ballistic regimes are then expected at each training step if the two-step sampling strategy is adopted. As we will show in Sec. 3.1, the two-step sampling method is more flexible than the first sampling approach in Sec. 2.2.1 for multiscale heat conduction problems. Theoretical analyses for providing intuition on the convergence of the two-step sampling approach is further present in A following [41].
@@ 3 Results and Discussion
In this section, we test the proposed method for simulating the heat conduction in solids ranging from diffusive to ballistic regimes, using four demonstration examples: (1) steady and unsteady quasi-one-dimensional (quasi-1D) film heat conduction, (2) quasi-two-dimensional (quasi-2D) heat conduction in a square domain, and (3) three-dimensional (3D) cuboid phonon transport. Further, the Adam optimizer, which is a variant of the stochastic gradient descent method, is utilized to minimize the loss functions in MC-PINNs. The details for computations, e.g., the architectures of NNs, activation functions, and the numbers of points used in Monte Carlo method/training steps, etc., as well as the details for reference solutions, are present in B.
@@ 3.1 Quasi-1D film heat conduction
We first consider a quasi-1D heat conduction problem across a film, as depicted in Fig. 4. The thickness of the film is $L=1$, which serves as the characteristic length here. The temperatures at the left and right boundaries, i.e., $x=0$ and $L$, are $T_{h}=1$ and $T_{c}=0$, respectively. We employ the MC-PINNs to model the heat conduction at four different Knudsen numbers, i.e., $\rm{Kn}=0.01,0.1,1$, and $10$, to justify the capability of the present method for multiscale heat conduction. For the quasi-1D problem, the angular variable can be expressed as $s_{x}=\cos{\theta}$. In all test cases, we randomly select $B_{\bm{x}}=40$ and $B_{\bm{s}}=16$ points in the spatial and angular domain for the computations of the loss of the equation at each training step, respectively. For the boundary conditions, we randomly select $B_{x_{w},\bm{s}}=50$ points in the solid angular space at the boundary $x=0$ and $x=1$ to implement the thermalization boundary condition in Eq. (6).
@@
To justify the accuracy of the present method, we first present the energy distribution functions computed from the MC-PINNs for the case with $\rm{Kn}=1$ at two representative locations, i.e., $x=0$ and $0.5$, in Fig. 5. As shown, the results from the MC-PINNs agree well with numerical results in [20]. In addition, the temperature as well as the heat flux are QoI in heat conduction problems rather than the energy distribution functions. We then illustrate the predicted temperature and the heat flux for $x\in[0,1]$ from MC-PINNs in Fig. 6. It is observed that: (1) all the results from the MC-PINNs are in good agreement with those in [17], (2) the temperature exhibits significant nonlinearities near the boundaries for the test cases with $\rm{Kn}>0.01$ (Fig. 6), and (3) the heat flux $q$ remains constant although the temperature is nonlinear for $\rm{Kn}>0.01$, which is clearly different from the widely used Fourier’s law (Fig. 6). In other words, the heat conduction for cases with $\rm{Kn}>0.01$ exhibits significant non-Fourier behavior.
@@
As shown, the relative $L_{2}$ errors between the present method and the reference solutions for all test cases are less than $1\%$, demonstrating that the MC-PINNs are capable of simulating heat conduction at both diffusive and ballistic scales with good accuracy. Additional results on the effect of $B_{\bm{s}}$, i.e., the number of points used at each training step in the angular domain, on the predicted accuracy are present in C.
Figure 6: Quasi-1D film heat conduction: Predicted (a) temperature, and (b) heat flux, from MC-PINNs at different $\rm{Kn}$. Symbols: Reference solution [17]; Solid line: MC-PINNs; Gray dashed line: the solution of Fourier’s Law.
Table 1: Quasi-1D film heat conduction: Relative $L_{2}$ errors between MC-PINNs and reference solution at different $\rm{Kn}$. $E_{\phi}$: Relative $L_{2}$ errors for temperature and heat flux.
@@
    PINNs-2: We employ $B_{\bm{x}}$ and $B_{\bm{s}}$ uniform discrete points in the spatial and solid angular domains, respectively. The tensor product is then utilized to generate the residual points at each training step, simialar as in PINNs-1.
@@
    PINNs-3: In the spatial domain, we use $B_{\bm{x}}$ uniform discrete points, and $B_{\bm{s}}$ quadrature points from the Gauss-Legendre method are employed in the solid angular domain, The residual points at each training step is then constructed using the tensor product. We note that this sampling approach is the same as in [35].
@@
    PINNs-4: We randomly sample $B_{\bm{s}}$ points in the solid angular domain, but use $B_{\bm{x}}$ uniform points in the spatial domain. The tensor product is then use to generate the residual points at each training step, simialar as in PINNs-1. Note that the sampling approach here can be viewed as a special case in PINNs-1 by replacing the random points with the fixed uniform grids in the spatial domain.
@@
    PINNs-5: Randomly sample $B_{\bm{x}}\times B_{\bm{s}}$ residual points in the spatial-angular domain at each training step. We note that this approach is one of the most widely used sampling methods in PINNs.
We utilize these five sampling methods in PINNs to model the heat conduction in the 1D film at $\rm{Kn}=0.01$, 1, and 10. The results and the computational errors are present in Fig. 7 and Table 2, respectively. As we can see, (1) the PINNs with the two-step sampling approach, i.e., PINNs-1 and PINNs-4, are able to provide the most accurate results for the heat conduction from diffusive to ballistic regimes; (2) the method in [35], i.e., PINNs-3, and PINNs-2, are able to achieve good accuracy for Kn $=0.01$ and 1, but failed to capture the non-Fourier’s behavior accurately at Kn $=10$ (e.g., the temperature jumps at the boundaries), as shown in Fig. 7; and (3) the computational errors for PINNs-5 increases with the decreasing Kn. As discussed in Sec. 2.2.2, here we only solve the BTE at one point in the solid angular space at each spatial point, leading to poor accuracy for the computation of $e^{eq}$. At the ballistic scale, i.e., cases with large Kn, we have very few collisions for the phonon, and the inaccurate estimation of $e^{eq}$ has little effect on the predicted accuracy. However, for cases with small Kn, we have frequent collisions of the phonon, and the inaccurate prediction of $e^{eq}$ results in poor accuracy for solving the BTE.
Figure 7: Quasi-1D film heat conduction: Effects of different sampling approaches in solid angular space on the predicted accuracy at (a) $\rm{Kn}=0.01$, (b) $\rm{Kn}=1$, and (c) $\rm{Kn}=10$. Symbols: Reference solution [18]; PINNs-1: Two-step sampling method: Random sampling in both spatial and solid angular domain; PINNs-2: uniform discrete points in both spatial and solid angular domain; PINNs-3: uniform discrete points in spatial domain and GL quadrature points in angular domain; PINNs-4: Two-step sampling method: uniform discrete points in spatial domain and random sampling in solid angular space; PINNs-5: random sampling in spatial-angular domain.
Table 2: Quasi-1D film heat conduction: Relative $L_{2}$ errors of the temperatures between PINNs and reference solution with different training points. PINNs-1 and PINNs-4: two-step sampling approach proposed in the current study; PINNs-2: uniform grids in both domains; PINNs-3: sampling approach used in [33, 35]; PINNs-5: sampling approach used in [34].
@@
$E_{T}(\rm{Kn}=0.01)$ 0.402% 0.383% 0.370% 0.390% 2.112%
$E_{T}(\rm{Kn}=1)$ 0.299% 1.005% 0.678% 0.214% 1.224%
$E_{T}(\rm{Kn}=10)$ 0.092% 1.280% 1.076% 0.065% 0.115%
3.2 Unsteady quasi-1D film heat conduction with space varying Knudsen number
In this section, we apply the MC-PINNs to model the time-dependent heat conduction in a quasi-1D film with a varying Knudsen number, which is defined as follows:
@@ 3.2 Unsteady quasi-1D film heat conduction with space varying Knudsen number
where $\rm{Kn}_{\max}$ and $\rm{Kn}_{\min}$ represent the maximum and minimum Knudsen number, respectively; $x_{c}=L/2$ is the center of the film with $L$ denoting the thickness of the film, and $d=0.01$ characterizes the thickness of the transition layer. In the present case, we set $\rm{Kn}_{\max}=10$ and $\rm{Kn}_{\min}=0.01$. The distribution of the Knudsen number within the film is shown in Fig. 8. It can be seen that we have a significant change in $\rm{Kn}$ near the center of the film. The heat conduction at the left half is at the ballistic regime, and is at the diffusive regime on the right half of the film.
@@
The computational domain as well as the boundary conditions are kept the same as in Sec. 3.1. In addition, we set $T(t=0,x)=(T_{h}+T_{c})/2$ in the entire domain. The time domain here is set as $t\in[0,10]$. In the computations, we sample $B_{t,x}=100$ random points in the temporal-spatial domain and $B_{s}=16$ random points in the solid angular domain. Further, $B_{t,x_{w},\bm{s}}=100$ random points are generated for boundary conditions and $B_{t_{0},x,\bm{s}}=100$ random points for initial conditions.
@@
We would like to point out that 100 Gauss-Legendre quadrature points are employed in the solid angular space in the numerical method to accurately capture the non-Fourier’s behavior [20]. In MC-PINNs, we only use 16 points in the solid angular domain at each training step, and it is capable of providing good results for heat conduction ranging from diffusive to ballistic scales. The above results indicate the great flexibility of the MC-PINNs. We do not need to carefully design the discretization of the angular domain, which is a long-standing challenge in conventional numerical methods for simulating multiscale heat conduction problems.
Figure 9: Unsteady quasi-1D film heat conduction: The predicted temperature at different times with space varying Knudsen number. (a) $t=0.2$, (b) $t=0.5$, (c) $t=1$, (d) $t=10$. Symbols: Reference solution [20]; Black solid line: MC-PINNs.
@@ 3.3 Quasi-2D square heat conduction
We now consider a quasi-2D heat conduction problem in a square domain with the length $L=1$. The temperature is set to $T_{h}=1$ at the top boundary and $T_{c}=0$ at the remaining boundaries, as shown in Fig. 10.
Figure 10: Schematic of the quasi-2D heat conduction in a square domain. $L$ is the characteristic length, $T_{h}$ is the upper plate dimensional temperature and $T_{c}$ is the dimensional temperature of the remaining plates.
In this section, we employ the MC-PINNs to model the heat conduction in the 2D square domain at $\rm{Kn}=0.1,1$ and $10$. In our computations, we first randomly sample $B_{\bm{x}}=100$ points in the spatial domain, and then select $B_{\bm{s}}=64$ random points in the solid angular domain $(s_{x},s_{y})$. As for boundary conditions, we randomly select $B_{\bm{x,s}}=100$ in the spatial-angular domain at each boundary to implement the thermalization boundary condition defined in Eq. (6).
The results of MC-PINNs are shown in Figs. 11 and 12. We can see that the predicted temperatures from the MC-PINNs are in good agreement with the reference solutions [18] in all test cases. We also observe in Figs. 11 and 12 there are significant temperature jumps at the left/right boundaries. In addition, the temperature jump becomes more pronounced with the increase of $\rm{Kn}$. This phenomenon is reasonable since the non-equilibrium effects become more pronounced in the transition and ballistic regimes, leading to distinct temperature discontinuities at the walls.
Figure 11: Quasi-2D heat conduction: Distribution of the temperature at (a) $\rm{Kn}=0.1$, (b) $\rm{Kn}=1$ and (c) $\rm{Kn}=10$. Colored background with black solid line: reference solution [18]; Red dashed line: the MC-PINNs.
Figure 12: Quasi-2D heat conduction: Temperature along the center vertical line at different $\rm{Kn}$. Symbols: Reference solution[18]; Solid line: MC-PINNs.
We further compute the relative $L_{2}$ error between the MC-PINNs and the reference solutions [18] at different $\rm{Kn}$ to demonstrate the accuracy of the present method in quasi-2D problems. As shown in Table 3, the relative $L_{2}$ errors for the three test cases are smaller than $2\%$, which again confirms the good accuracy of MC-PINNs for modeling multiscale heat conduction.
Table 3: Quasi-2D heat conduction problem: Relative $L_{2}$ errors between MC-PINNs and reference solution at different $\rm{Kn}$.
- $\rm{Kn}=0.1$ $\rm{Kn}=1$ $\rm{Kn}=10$
@@
In this section, we proceed to consider the phonon transport in a 3D cuboid. The temperature is set as $T_{h}$ at the top boundary and $T_{c}$ at the remaining boundaries, as shown in Fig. 13. Specifically, we set $T_{h}=1$, $T_{c}=0$, and $L=1$.
@@
In the MC-PINNs, the two-step sampling strategy is applied for generating the training points for the loss of residual with $B_{\bm{x}}=100$ in the spatial domain and $B_{\bm{s}}=64$ in the solid angular domain $(s_{x},s_{y},s_{z})$. For the boundary conditions, we randomly select $B_{\bm{x,s}}=100$ in the spatial-angular domain for each wall at each training step. The boundary condition implemented here is defined in Eq. (6), which is the same as the previous two cases in Secs. 3.1 and 3.3. As for the reference solution, we employ the implicit kinetic scheme to solve the stationary phonon BTE developed in [18]. We note that the numerical method in [18] employs a macroscopic equation to accelerate convergence in the diffusive regime. Additionally, it incorporates a memory reduction technique to reduce memory usage in computations. We note that $40\times 40$ and $80\times 80$ discrete points in the angular space are employed in the numerical method for $\rm{Kn}=0.1,1$ and $\rm{Kn}=10$, respectively, to ensure the convergence. More details for the discretization of the spatial space in the numerical method are present in B.
The predicted temperature from MC-PINNs in each case at $x=0.5$ are present in Figs. 14 and 15. Again, good agreements are observed for all test cases between the MC-PINNs and the reference solutions. Similar to in the 2D cases, there are significant temperature jumps at the boundaries when $\rm{Kn}\geq 1$, which have been discussed in Sec. 3.3 and will not present here for simplicity. We then illustrate the relative $L_{2}$ errors between the temperature from the MC-PINNs and the reference solutions at different $\rm{Kn}$ in Table 4. The results further confirm the capability of the MC-PINNs for 3D multiscale heat conduction problems.
@@
Figure 15: 3D heat conduction problem: Predicted temperature from MC-PINNs at $x=0.5$ and $y=0.5$. Symbols: Reference solution [18]; Solid line: MC-PINNs.